# On the Feasibility of Hijacking MLLMs’ Decision Chain via One Perturbation

### v1

@@ Appendix A Design Details of RIST Dataset
We introduce the Real Image Sequence Trajectories (RIST), a newly annotated dataset to evaluate adversary capacity to induce distinct, predefined targets by exploiting the fine-grained semantics. The fine-grained semantics are defined by a sequence of images with similar content, such as images of a specific intersection. This grants the adversary “switch-case” capability to “program” the incidents flexibly. For instance, a single perturbation could first guide an MLLM to output “grasp” when a robot observes a knife, and then trigger a dangerous action like “throw” when a human subsequently enters the scene.
@@ A.1 Dataset Curation and Content
AutoDriving is designed to simulate the decision chain of a driving autonomous vehicle. It comprises 15 trajectories, each further segmented into 5 distinct categories. Each category is represented by a sequence of 10 POV images, formatted as 1920x1080 pixel JPG images. The initial POV driving videos for AutoDriving are sourced from the OpenDV-2K dataset [38] and the ONCE dataset [25]. This curation spans multiple countries and varies in temporal and meteorological conditions to ensure diversity.
RoboTasking is curated to evaluate the robustness of robotic decision chains in realistic indoor environments. It consists of 13 trajectories, each of which is divided into 2 categories. Similarly, each category comprises a sequence of 10 POV images, captured at a resolution of 320x180 pixels and stored as JPG files. The original POV robot videos for RoboTasking are sourced from the DROID dataset [33], utilizing the robot’s wrist-mounted camera perspective. We cover a wide range of robotic tasks (e.g., grasping objects, placing items, and operating tools) across various environments, including kitchens, living rooms, and workshops.
@@ A.2 Target Action Annotation
For the AutoDriving, the image sequence corresponding to each category is input into Gemini-2.5-pro [8]. The model is prompted to act as an autonomous driving system and select a driving action based on the input images that would most likely pose a significant traffic safety risk. This selected action is assigned as the target for this category.
@@
Appendix B Optimization Algorithm
This section presents a description of the optimization strategy for generating SAUPs, as illustrated in Algorithm 1.
Given an image dataset $\mathcal{D}$, the adversary assigns a target label to each image based on its category. Thus, we can reorganize $\mathcal{D}$ such that each data entry consists of an image $\mathbf{x}$, a corresponding target label $\mathbf{{t}}$, and a query prompt $\mathbf{p}$.
For normalized space optimization, we first apply the image preprocessing function $f_{\mathcal{T}}$ to both the dataset $\mathcal{D}$ and the initialized perturbation $\delta$ in lines 2. Each image $\mathbf{x}$ in the dataset is transformed into a representation $\mathbf{Z}$, while the perturbation $\delta$ is transformed into $\Delta$. To enable continuous optimization in the normalized space, we further transform the image bounds and perturbation constraints accordingly. In lines 3 of the algorithm, the image pixel value bounds, originally in the range $\left[0,1\right]$, are transformed to the corresponding bounds $\left[\underline{\mathbf{Z}},\overline{\mathbf{Z}}\right]$ in the normalized space. Simultaneously, the perturbation constraints are transformed into new region constraints $\mathbf{R}$.
For semantic separation optimization, we begin each optimization iteration by clipping the perturbed image with the new constraint, as shown in Line 6. These inputs are then fed into the dual-loss function $\mathcal{L}$, which integrates cross-entropy loss and margin loss. This objective not only enhances alignment between inputs and their assigned targets but also encourages separation among different target classes. Finally, the perturbation is updated via gradient descent using the learning rate $\eta$.
After completing the maximum number of optimization epochs, we manually construct an inverse transformation $f^{-1}_{\mathcal{T}}$ to map the optimized perturbation back to the original pixel space. The final perturbation $\delta$ can be generalized to unseen images, support multiple targets via semantic-aware, and adhere to perturbation constraints.
Input: Dataset $\mathcal{D}$, each entry consisting of an image $\mathbf{x}$, its corresponding target label $\mathbf{t}$, and user prompt $\mathbf{p}$. Semantic separation loss function $\mathcal{L}$. Image preprocess function $f_{\mathcal{T}}$. Perturbation constraint $\mathbf{r}$. Learning rate $\eta$.
Output: A semantic-aware universal perturbation $\delta$.
1
21exRandomly initialize $\delta$ ;
3 $\mathcal{D}^{\prime},\Delta\leftarrow f_{\mathcal{T}}\left(D\right),f_{\mathcal{T}}\left(\delta\right)$ ;
4 $\underline{\mathbf{Z}},\overline{\mathbf{Z}},\mathbf{R}\leftarrow f_{\mathcal{T}}\left(\mathbf{0}\right),f_{\mathcal{T}}\left(\mathbf{1}\right),f_{\mathcal{T}}\left(\mathbf{R}\right)$ ;
5
6for $i\leftarrow 1$ to EpochEpoch do
7
8 foreach $\left\{\mathbf{Z},\mathbf{t},\mathbf{p}\right\}\in\mathcal{D}^{\prime}$ do
9    $\mathbf{Z}^{\prime}\leftarrow\left[\mathbf{Z}\cdot(1-\mathbf{R})+\mathbf{{R}}\cdot\Delta\right]_{\underline{\mathbf{Z}}}^{\overline{\mathbf{Z}}}$ ;
10    $L\leftarrow\mathcal{L}\left(\mathbf{Z}^{\prime},\mathbf{t;\mathbf{p}}\right)$ ;
11    $\Delta\leftarrow\Delta+\eta\cdot\nabla L$ ;
12
13   end foreach
14
15 end for
16Construct inverse preprocess function $f^{-1}_{\mathcal{T}}$;
17 $\delta\leftarrow f^{-1}_{\mathcal{T}}\left(\Delta\right)$ ;
18 return $\delta$
Algorithm 1 Optimization Algorithm for SAUPs
Appendix C Evaluation on Multiple Prompts.
@@ 17 $\delta\leftarrow f^{-1}_{\mathcal{T}}\left(\Delta\right)$ ;
Appendix D Impact of Perturbation Region
@@
Table 4: Attack success rates under different sizes of training set
Model Constraint Size of Training Set
10 20 30 40 50
Llava Frame Train $1.00$ $1.00$ $0.93$ $0.93$ $0.96$
Test $0.10$ $0.53$ $0.58$ $0.83$ $0.95$
Corner Train $1.00$ $1.00$ $0.98$ $0.94$ $0.90$
Test $0.45$ $0.75$ $0.88$ $0.83$ $0.88$
Qwen Frame Train $1.00$ $0.98$ $0.92$ $0.96$ $0.95$
Test $0.85$ $0.63$ $0.85$ $0.95$ $0.93$
Corner Train $1.00$ $1.00$ $1.00$ $1.00$ $1.00$
Test $0.60$ $0.73$ $0.80$ $0.83$ $0.98$
Intern Frame Train $1.00$ $1.00$ $1.00$ $1.00$ $1.00$
Test $0.65$ $0.98$ $0.90$ $1.00$ $0.98$
Corner Train $1.00$ $1.00$ $1.00$ $1.00$ $1.00$
Test $0.65$ $0.90$ $0.85$ $1.00$ $1.00$
Appendix E Impact of Training Set Size
@@
Figure 10: Illustration of confidence per tokens of each target, where Tii denotes the start of the $i$-th target.
Appendix F Analysis on Target Token Confidence
We analyze the confidence scores of perturbed images for five different targets. Since each target consists of one or more tokens, this section further decomposes each target into its tokens to display the confidence score of each token. As illustrated in Figure 10, we observe that perturbed images yield high confidence scores across multiple target tokens. For instance, Perturbed Image Cluster 1 displays elevated confidence for both token 2 and token 3 of Target 2. These results underscore the necessity of incorporating a semantic segmentation module in our design.
Appendix G Discussion and Limitation
G.1 Discussion
@@
Current research on adversarial perturbations encounters limitations in flexibility, which can be categorized as shown in Table 5. Traditional adversarial perturbations [20, 35] are typically optimized on a per-instance basis. These methods can be either untargeted, aiming only to cause misclassification, or single-targeted, where the goal is to force the model to produce a specific incorrect label. However, these methods fail to scale or generalize to broader scenarios due to their inflexible, instance-specific design. Universal adversarial perturbations (UAPs) [26, 27] extend adversarial attacks by creating a single perturbation applicable across multiple inputs. Untargeted UAPs are effective in degrading model accuracy broadly, while single-targeted UAPs attempt to shift model predictions toward a fixed label across all inputs. Despite their improved generalization, the perturbations fail to provide fine-grained control over the output based on the semantics of the input.
For seeking to manipulate multiple decisions towards distinct targets, existing approaches demand an individual perturbation for each input. This per-input generation makes them impractical for real-world applications, such as physically applying perturbations to every frame of a video.
In this context, SAUPs enable the adversary to freely map multiple images to various targets, representing a more sophisticated “switch-case” attack capability.
G.2 Limitation
While this paper demonstrates the existence and potential threat of SAUPs, it also has some limitations.
White-box Assumption and Transferability. This paper is based on the white-box assumption, where the attacker has full access to the target model’s architecture and parameters. While this assumption is standard in current image hijack research [4], it poses a significant limitation in real-world scenarios, where models are often black boxes.
For cross-model transferability, prior work [29] has demonstrated that MLLMs are resistant to transfer attacks, and current studies on image hijacking [4, 24, 23] likewise have not observed cross-model transferability. Given that SAUPs impose even more stringent spatial constraints, we also do not observe transferability of SAUPs across the three MLLMs considered in this study.
Robustness and Physical Application. Although we demonstrate the feasibility of hijacking the decision chain via one perturbation, applying such a perturbation to the physical environment remains a significant challenge. The robustness of these perturbations to physical factors, such as lighting conditions and motion blur, requires further study.

### v2

@@ Appendix A Design Details of RIST Dataset
We introduce the Real Image Sequence Trajectories (RIST), a newly annotated dataset to evaluate an adversary’s capacity to induce distinct, predefined targets by exploiting fine-grained semantics. The fine-grained semantics are defined by a sequence of images with similar content, such as images of a specific intersection. This grants the adversary “switch-case” capability to “program” the incidents flexibly. For instance, a single perturbation could first guide an MLLM to output “grasp” when a robot observes a knife, and then trigger a dangerous action like “throw” when a human subsequently enters the scene.
Semantic classification within a continuous trajectory is not supported by common image classification (Deng et al., 2009) or image captioning datasets. COCO Captions (Chen et al., 2015) and Flickr30k Entities (Plummer et al., 2015) treat each image as an independent sample, providing multiple annotated captions per image. While region-level context is occasionally included, these benchmarks fail to maintain temporal continuity or a consistent perspective across samples.
RIST distinguishes itself from existing temporal datasets, which often lack classification based on semantic similarity within a single trajectory. While MSR-VTT (Xu et al., 2016) provides clip-level captions and temporal context, it does not define semantic categories within trajectories. Similarly, autonomous driving datasets like KITTI (Geiger et al., 2013), nuScenes (Caesar et al., 2020), and Waymo Open Dataset (Sun et al., 2020) offer continuous sequences for perception and prediction tasks; however, they are not organized by event-centric categories, nor do they feature predefined goals for each category. In contrast, RIST segments each consistent-viewpoint trajectory into multiple semantic categories and assigns a specific predefined goal to each.
Figure 11: Illustration of RIST, which consists of real-world image trajectories across two scenarios: AutoDriving and RoboTasking.
@@ A.1 Dataset Curation and Content
AutoDriving is designed to simulate stateless decisions of a driving autonomous vehicle. It comprises 15 trajectories, each further segmented into 5 distinct categories. Each category is represented by a sequence of 10 POV images, formatted as 1920x1080 pixel JPG images. The initial POV driving videos for AutoDriving are sourced from the OpenDV-2K dataset (Yang et al., 2024) and the ONCE dataset (Mao et al., 2021). This curation spans multiple countries and varies in temporal and meteorological conditions to ensure diversity.
RoboTasking is curated to evaluate the robustness of robotic stateless decisions in realistic indoor environments. It consists of 13 trajectories, each of which is divided into 2 categories. Similarly, each category comprises a sequence of 10 POV images, captured at a resolution of 320x180 pixels and stored as JPG files. The original POV robot videos for RoboTasking are sourced from the DROID dataset (Tang et al., 2025), utilizing the robot’s wrist-mounted camera perspective. We cover a wide range of robotic tasks (e.g., grasping objects, placing items, and operating tools) across various environments, including kitchens, living rooms, and workshops.
@@ A.2 Target Action Annotation
For the AutoDriving, the image sequence corresponding to each category is input into Gemini-2.5-pro (DeepMind, 2025). The model is prompted to act as an autonomous driving system and select a driving action based on the input images that would most likely pose a significant traffic safety risk. This selected action is assigned as the target for this category.
@@
Appendix B Evaluation on Multiple Prompts.
@@
Appendix C Impact of Perturbation Region
@@
Algorithm 1 SORT Optimization
0: Training set $\mathcal{D}=\{(x^{(c)},t^{(c)})\}$ containing images and target prompts;
0: Loss Function $\mathcal{L}_{Total}$, Normalization function $\Psi$
0: Learning rate $\eta$, Perturbation Mask $M$
0: Semantic-Aware Universal Perturbation $\delta$
1:Initialize: $\Delta\leftarrow\mathbf{0}$ {Initialize variable in normalized space}
2:Initialize: $\Delta_{L}$, $\Delta_{U}\leftarrow\Psi(\mathbf{0}),\Psi(\mathbf{1})$ {Initialize lower and upper bound in normalized space}
3:while not converged do
4:  Sample batch of images and targets $(x,t)$ from $\mathcal{D}$
5:  $x_{adv}=\Psi(x)+\Delta$ {Construct adversarial images}
6:  $Loss\leftarrow\mathcal{L}_{Total}(x_{adv},t)$ {Compute loss}
7:  $\Delta\leftarrow\Delta-\eta\cdot\nabla_{\Delta}Loss$ {Update variable}
8:  $\Delta\leftarrow\text{clip}(\Delta\odot M,\Delta_{L},\Delta_{U})$ {Constraint perturbation}
9:end while
10:Return $\delta=\Psi^{-1}(\Delta)$
Appendix D Algorithm Description
In this section, we provide the detailed pseudocode in Algorithm 1 for the SORT optimization. To ensure optimization stability, SORT performs a change of variables from the pixel space to the normalized space. While mathematically simple, this reparameterization proves highly effective in mitigating gradient instability and avoiding local optima. Instead of optimizing the perturbation $\delta$ directly, we define a trainable variable $\Delta$ in the normalized feature space, constructing the adversarial input as $x_{adv}=\Psi(x)+\Delta$.
Since $\Delta$ should be converted to the valid pixel range $[0,1]$ through inverse transformation $\Psi^{-1}$ at the end of the algorithm, an unconstrained $\Delta$ will be significantly damaged due to truncation. To prevent this, we clip $\Delta$ within the bounds $\Psi(\mathbf{0}),\Psi(\mathbf{1})]$ at each iteration. This mechanism guarantees that every update to the perturbation in the normalized space translates accurately to the physical image space.
Appendix E Impact of the Size of Training Set
@@
Table 4: Attack success rates under different sizes of training set
Model Frame Corner
10 20 30 40 50 10 20 30 40 50
Llava Train $1.00$ $1.00$ $0.93$ $0.93$ $0.96$ $1.00$ $1.00$ $0.98$ $0.94$ $0.90$
Test $0.10$ $0.53$ $0.58$ $0.83$ $0.95$ $0.45$ $0.75$ $0.88$ $0.83$ $0.88$
Qwen Train $1.00$ $0.98$ $0.92$ $0.96$ $0.95$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$
Test $0.85$ $0.63$ $0.85$ $0.95$ $0.93$ $0.60$ $0.73$ $0.80$ $0.83$ $0.98$
Intern Train $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$
Test $0.65$ $0.98$ $0.90$ $1.00$ $0.98$ $0.65$ $0.90$ $0.85$ $1.00$ $1.00$
@@
Appendix F Discussion
@@
Current research on adversarial perturbations encounters limitations in flexibility, which can be categorized as shown in Table 5. Traditional adversarial perturbations (Lin et al., 2019; Wang and He, 2021) are typically optimized on a per-instance basis. These methods can be either untargeted, aiming only to cause misclassification, or single-targeted, where the goal is to force the model to produce a specific incorrect label. However, these methods fail to scale or generalize to broader scenarios due to their inflexible, instance-specific design. Universal adversarial perturbations (UAPs) (Moosavi-Dezfooli et al., 2017; Naseer et al., 2021) extend adversarial attacks by creating a single perturbation applicable across multiple inputs. Untargeted UAPs are effective in degrading model accuracy broadly, while single-targeted UAPs attempt to shift model predictions toward a fixed label across all inputs. Despite their improved generalization, the perturbations fail to provide fine-grained control over the output based on the semantics of the input.
For seeking to manipulate multiple decisions towards distinct targets, existing approaches demand an individual perturbation for each input. In contrast, SAUPs enable the adversary to freely map multiple images to various targets, representing a more sophisticated “switch-case” attack capability.